OpenStudy (anonymous):

solve the following problems the difference between six times a number and 9 equal to five times the sum of the number 12. find the number

OpenStudy (anonymous):

five times the sum of the number 12??

OpenStudy (anonymous):

no idea how to even start this

OpenStudy (anonymous):

u have written in d question....equal to five times the sum of number 12. ...which is not clear to me.

OpenStudy (anonymous):

This is how it goes \[(6k-9)=12\times5\] Which is 6k-9=60 6k =69 and so k = 69/6

OpenStudy (anonymous):

@akash123 If I say that instead of adding a number n times I multiply it with n, it is the same thing. You are confusing urself with the word number...

OpenStudy (anonymous):

so what should i do

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

@akash123 See that It is only taking about the number 12 in the question and the rest is to confuse you It is just trying to say that If a add 12+12+12+12+12 this is equal to 12x5

OpenStudy (anonymous):

ok...sorry

OpenStudy (anonymous):

And @teresaa Just try to break the question into two parts first. The first part says "between six times a number and 9" Do it This will be 6k-9 Now the second part of the question This says "five times the sum of the number 12" Here just remove the word the from the sentence and you will understand what I want to say i.e five times....sum of number 12 Fine it and the equate the two parts of the question